To estimate animal population, biologists count the total number of animals in a small section of a habitat. The total population of animals is directly proportional to the size of the habitat(in acres) polled.
a.Write an equation using only one variable that could be used to solve for the constat of variation k.
b.A biologist counted 12 white tail deer in a 100 acre parcel of land in a nature preserve.Find the constant of variation k.
c.If the entire nature preserve is 2,500 acres, then what is the total white tail deer population in the preserve? Describe the answer?Thanks Answer by stanbon(75887) 
 To estimate animal population, biologists count the total number of animals in a small section of a habitat. The total population of animals is directly proportional to the size of the habitat(in acres) polled.
----
a.Write an equation using only one variable that could be used to solve for the constat of variation k.
---
T = ka
------
b.A biologist counted 12 white tail deer in a 100 acre parcel of land in a nature preserve.Find the constant of variation k.
12 = k*100
k = 12/100 = 0.12
--------------------------
c.If the entire nature preserve is 2,500 acres, then what is the total white tail deer population in the preserve? Describe the answer?Thanks
---
T = 0.12*2500
T = 300 deer
=============================
Cheers,
Stan H.
=============================